Two angles of a triangle measure 20 and 30 degrees.The third angle measures how many degrees?
Answer by 2897696(96)

First you must know that the total degree of an triangle, which is 180 degree.
then we use the information given to make an equation.
Info: let x= unknown angle, x#2=20, and x#3=30
MAKING EQUATION
Since we know that the total degree of a triangle is 180, we have to add all the other 3 angles together, to get 180.
EQUATION: 20+30+x=180
solve for x
50+x=180 (added 20+30)
x=130 (180-50)
Answer: the third degree of the triangle is 130 degrees.